Grande03:

Complete the sequence for deriving the quadratic formula using the quadratic equation in standard form, ax²+bx+c = 0

Grande03:

AZ:

I wouldn't really expect anyone to remember how to derive the quadratic formula You have ax^2 + bx + c = 0 The first thing you want to do is divide everything by 'a' so that way you can get the equation to be x^2 + ... = 0 After that, you want to move the third term to the other side Then you complete the square Once you complete the square, you'll get it in the form of (x + ??)^2 = something else and then you're just solving for 'x' at this point, you take the square root on both sides and then just get 'x' all by itself

AZ:

I'm not sure when you'll be back online so that I can walk you through each step and I doubt my explanation would be of much use (especially since it's a derivation of a formula and it's not a simple concept to grasp adding on to the fact that we're just taught to memorize the formulas and not derive them) So I did find this webpage that explains it in a more understandable way https://www.mathsisfun.com/algebra/quadratic-equation-derivation.html

Creed:

Grande, I would help you, you know I would. But, I don't understand how this is done. I knew how to do the question from yesterday. Sorry I couldn't help. ₵ⱤɆɆĐ

Grande03:

I actually got it right I need to close this lol sorry

Grande03:

AZ:

Good job! :)